# Pulse shape simulation for the reduced charge collection layer in $p$-type high-purity germanium detectors

### v1

@@ 1 Introduction
$P$-type high-purity germanium (HPGe) detectors have been a mainstay in nuclear and particle physics for decades due to their excellent energy resolution. Applications that require a detailed understanding of the detector signal response still face challenges though: experiments like rare event searches can take advantage of the intrinsically low background of HPGe. However, the intrinsic background of the detector and especially the surrounding parts is still not low enough for ultralow-background experiments such as direct dark matter detections zhao_first_2013 ; jiang_limits_2018 ; ma_results_2020 ; li_limits_2013 , neutrinoless double-beta decay searches dai_search_2022 ; zhang_searching_2024 ; abgrall_large_2017 , and coherent elastic neutrino-nucleus scattering detections wong_research_2006 ; bonet_constraints_2021 ; yang_recode_2024 ; belov_new_2025 . Such experiments use detailed analysis of the detector signal shapes to distinguish signal from background events to further lower their effective background, but this is not an easy task.
The inhomogeneous electrical and weighting fields of large-volume HPGe detectors, together with the distinct drift behavior between bulk and surface regions, are both a blessing and a curse for such use cases: they result in different signal shapes for different event topologies that are associated with signal and background. However, they also make faithful simulation of such detectors a hard problem. Because of this, experiments have often relied on a data-driven approach to build digital pulse shape processing (DSP) procedures and statistical and machine-learning models for background discrimination zhang_machine_2024 ; herrero-garcia_semi-supervised_2022 ; agostini_pulse_2022 ; holl_deep_2019 ; delaquis_deep_2018 ; radovic_machine_2018 ; cohen_machine_2018 . This data-driven approach has been successful, but is also limited, as the underlying truth of the data sets that the models are built on is not known event by event. Simulation-based and data-verified models would be the gold standard and have the potential for higher discriminatory power with better control of systematic effects.
@@
For $p$-type HPGe detectors in low-background experiments, the surface plays a special role. The surface $n^{+}$ electrode of a $p$-type HPGe detector is typically fabricated through thermal diffusion of lithium, resulting in a heavily doped region with a thickness of ~ 0.5–1 mm. However, this process introduces a layer in the $n^{+}$ electrode that is almost inactive close to the surface, with its charge collection efficiency increasing as a function of depth: due to the weak electric field in this region and a higher trapping (or recombination) rate dai_modeling_2023 , charge carriers are prone to sudden termination, leading to incomplete charge collection. Thus, charge depositions in this layer are either not registered at all or exhibit a measured energy lower than their true energy. It is sometimes called dead layer, inactive layer, or transition layer in the literature. For clarity, we will call it the reduced charge collection (RCC) layer in this paper. These RCC layer events constitute a significant background component in rare-event searches li_differentiation_2014 ; yang_bulk_2018 ; li_identification_2022 ; zhang_machine_2024 ; wang_bulk-surface_2025 .
@@
We want to study the charge drift of RCC layer events and develop a pulse shape simulation method that can both deepen our understanding of the underlying detector physics and enable a more precise evaluation of the signal efficiencies of pulse shape discrimination techniques dai_search_2022 ; zhang_searching_2024 . Realistic simulations can also generate clean datasets with known ground truth for developing advanced discrimination techniques zhang_machine_2024 .
@@ 2 Method
With "sensitive region", we refer to the detector bulk volume with full charge collection efficiency (CCE). RCC layer refers to the outside, including the whole $n$-type neutral region and a small part of the $n$-type depleted region, as shown in Figure 1. Following the convention of our previous research dai_modeling_2023 , the depth from the detector surface to the boundary between the sensitive region and the RCC layer is termed the full charge collection depth (FCCD), which is also conventionally regarded as the depth of the RCC layer.
@@ 2.1 Ionized impurity density profile
Ionized impurities refer to impurities with ionized atoms regardless of whether the carriers are depleted, including donors (doped) and acceptors (already existing in $p$-type Ge crystal). To model the physical processes in the RCC layer, we first need to establish the ionized impurity density profiles. For acceptor impurities, we assume a constant density dai_modeling_2023 . For donor impurities, which are introduced through thermal diffusion at the $n^{+}$ electrode, this work follows the calculations described in Reference dai_modeling_2023 . Therefore, the net ionized impurity density profile throughout the crystal is described by
$N_{\text{I,net}}(x)=N_{\text{d}}(x)-N_{\text{a}}\,$, (1)
$N_{\text{d}}(x)=N_{\text{s}}\cdot\text{erfc}(\frac{x}{2\sqrt{D_{\text{Li}}t_{\text{an}}}})\,$, (2)
$\log_{10}{(N_{s}/\text{cm}^{-3})}=21.27-(\frac{2610\ \text{K}}{T_{\text{an}}})\,$, (3)
$D_{\text{Li}}=D_{\text{0}}\cdot\exp(-\frac{H}{RT_{\text{an}}})\,$, (4)
@@ 2.3 Charge-carrier diffusion, self-repulsion, and trapping
We employ a constant-lifetime trapping model dai_modeling_2023 ; mei_impact_2016 ; mei_impact_2017 ; boggs_numerical_2023 , but assume different lifetimes in the sensitive region and the RCC layer. Therefore, our trapping model has four parameters: electron/hole lifetimes in the RCC layer $\tau_{\text{e/h,R}}$, and in the sensitive region $\tau_{\text{e/h,S}}$, respectively. For this paper, we set $\tau_{\text{e/h,S}}$ to 1 ms knoll_radiation_2010 , a typical value for high-purity germanium. However, in the RCC layer, the extremely high donor impurity density makes carriers more prone to trapping, resulting in significantly shorter lifetimes than those in the sensitive region. The lifetimes $\tau_{\text{e/h,R}}$ must be determined by experimental calibration. According to the Shockley-Ramo theorem he_review_2001 , the induced signal $Q_{\text{e/h}}(t)$ is given by
$Q_{\text{e/h}}(t)$ $=Q_{\text{e/h, A}}(t)+Q_{\text{e/h, R}}(t)+Q_{\text{e/h, S}}(t)\,$, (15)
$Q_{\text{e/h, A}}(t)$ $=w(\widehat{r}_{\text{e/h}}(t))q_{\text{e/h}}(t)\,$, (16)
$Q_{\text{e/h, TR}}(t)$ $=\sum_{i\in R(t)}{w(\widehat{r}_{\text{e/h}}(t_{i}))q_{\text{e/h}}(t_{i})\frac{dt_{i}}{\tau_{\text{e/h, R}}}}\,$, (17)
$Q_{\text{e/h, TS}}(t)$ $=\sum_{i\in S(t)}{w(\widehat{r}_{\text{e/h}}(t_{i}))q_{\text{e/h}}(t_{i})\frac{dt_{i}}{\tau_{\text{e/h, S}}}}\,$, (18)
$q_{\text{e/h}}(t)$ $=q_{\text{e/h}}(0)-q_{\text{e/h, R}}(t)-q_{\text{e/h, TS}}(t)\,$, (19)
$q_{\text{e/h, TR}}(t)$ $=\sum_{i\in R(t)}q_{\text{e/h}}(t_{i-1})\frac{dt_{i}}{\tau_{\text{e/h, R}}}\,$, (20)
$q_{\text{e/h, TS}}(t)$ $=\sum_{i\in S(t)}q_{\text{e/h}}(t_{i-1})\frac{dt_{i}}{\tau_{\text{e/h, S}}}\,$, (21)
where $dt_{i}$ is the time step length for the $i$-th step of drift, $Q_{\text{e/h, A}}(t)$ denotes the induced signal contributed by alive carriers at time $t$, $Q_{\text{e/h, TR}}(t)$ denotes the induced signal contributed by carriers trapped in the RCC layer, $Q_{\text{e/h, TS}}(t)$ denotes the induced signal contributed by carriers trapped in the sensitive region, $q_{\text{e/h}}(t)$ denote the charges of alive carriers, $q_{\text{e/h, TR}}(t)$ denote the charges of carriers trapped in the RCC layer, $q_{\text{e/h, TS}}(t)$ denote the charges of carriers trapped in the sensitive region, $\widehat{r}_{\text{e/h}}(t)$ are the carrier drift paths, $w(\widehat{r}_{\text{e/h}}(t))$ are the weighting potentials, $R(t)/S(t)$ denote the parts of the drift path in the RCC layer and the sensitive region, respectively.
@@ 2.4 Charge collection efficiency
Based on the aforementioned physical models, we can simulate both the paths of carriers and the induced signals of RCC layer events. Furthermore, we can calculate the CCE curve of the RCC layer $f_{\text{CCE}}$ and its uncertainty dai_modeling_2023 ; ma_study_2017 with
$f_{\text{CCE}}(x)$ $=\frac{A_{\text{p}}(x)}{Ne}\,$, (22)
$\frac{\sigma_{f}}{f}$ $=\frac{1}{\sqrt{N}}=\frac{1}{\sqrt{E/{E_{0}}}}\,$, (23)
where $x$ is the depth of the simulated single-site event from the detector surface, $A_{\text{p}}$ is the amplitude of the simulated charge pulse, $N$ is the initial number of the simulated charge carrier pairs, and $E_{0}$ is the ionization energy for germanium.
The CCE curve has been proven to be crucial for interpreting experimental spectra and correcting distortions in the raw simulated energy spectra in the low-energy region because alpha, beta, and low-energy gamma particles from outside have a higher probability to deposit energy close to the surface dai_modeling_2023 ; ma_study_2017 . For each event, we denote the simulated deposited energies and depths with $(E_{i},x_{i})$. The raw simulated total energy $E_{\text{raw}}$, the total energy corrected with CCE curve $E_{\text{w/}}$, and total energy corrected without CCE curve $E_{\text{w/o}}$ are given by
@@
Figure 4: Illustration of the CCE curve $f_{\text{CCE}}(x)$ and the step function $\Theta(x-\text{FCCD})$ with FCCD = 0.9 mm.
For $p$-type HPGe detectors, in most cases, electrons in the RCC layer contribute negligibly to the induced signal on the $p^{+}$ contact because the weighting potential there is nearly zero li_identification_2022 ; dai_pulse_2024 ; zhang_searching_2023 . Thus, the CCE approximately equals the fraction of holes that escape the RCC layer before getting trapped or recombined, reach the sensitive region, and get collected by the $p^{+}$ electrode. One can determine the FCCD through pulse shape simulation but also experimentally, by matching the experimentally measured energy spectrum with the energy spectrum simulated using a Monte-Carlo software like Geant4 jiang_measurement_2016 ; ma_study_2017 ; agostinelli_geant4simulation_2003 ; allison_geant4_2006 ; allison_recent_2016 .
@@ 3 Method validation
In the theory-based approach, we compute the CCE in the RCC layer by solving the one-dimensional hole transport equation analytically. For this, we consider three holes-transport physics processes: drift under the electric field, diffusion, and trapping.
@@ 3.2 Charge collection efficiency validation
Figure 7 shows the distribution of the time when the holes arrive at the $p$-$n$ boundary for a starting depth of 0.5 mm from the surface. The integral of this curve divided by the initial hole number (3389) is the charge collection efficiency of the event, see Equation 22. We observe that the analytical results lie within 1 $\sigma$ of the simulations (using the estimated Poisson fluctuation).
@@
In addition to the arrival time, we compare the depth dependency of the CCE by generating events at different depths in the RCC layer. Figure 8 shows that the two CCE curves are also in very good agreement between numerical simulation and analytical solution.

### v2

@@ 1 Introduction
$P$-type high-purity germanium (HPGe) detectors have been a mainstay in nuclear and particle physics for decades due to their excellent energy resolution. Applications that require a detailed understanding of the detector signal response still face challenges though: experiments like rare event searches can take advantage of the intrinsically low background of HPGe detectors. However, the intrinsic background of the detector and especially the surrounding parts is still not low enough for ultralow-background experiments such as direct dark matter detections zhao_first_2013 ; jiang_limits_2018 ; ma_results_2020 ; li_limits_2013 , neutrinoless double-beta decay searches dai_search_2022 ; zhang_searching_2024 ; abgrall_large_2017 , and coherent elastic neutrino-nucleus scattering detections wong_research_2006 ; bonet_constraints_2021 ; yang_recode_2024 ; belov_new_2025 . Such experiments use detailed analysis of the detector signal shapes to distinguish signal from background events to further lower their effective background, but this is not an easy task.
The inhomogeneous electrical and weighting fields of large-volume HPGe detectors, together with the distinct drift behavior between bulk and surface regions, are both a blessing and a curse for such use cases: they result in different signal shapes for different event topologies that are associated with signal and background martin_determining_2012 ; agostini_pulse_2022 ; comellato_topologies_2023 ; dai_virtual_2024 . However, they also make faithful simulation of such detectors a hard problem. Because of this, experiments have often relied on a data-driven approach to build digital pulse shape processing (DSP) procedures and statistical and machine-learning models for background discrimination zhang_machine_2024 ; agostini_pulse_2022 ; holl_deep_2019 ; aalseth_maximum_2015 ; abt_alpha-event_2017 ; arnquist_alpha_2022 ; dai_virtual_2024 ; bonet_pulse_2024 ; radovic_machine_2018 ; cohen_machine_2018 . This data-driven approach has been successful, but is also limited, as the underlying truth of the data sets that the models are built on is not known event by event. Simulation-based and data-verified models would be the gold standard and have the potential for higher discriminatory power with better control of systematic effects.
@@
For $p$-type HPGe detectors in low-background experiments, the surface plays a special role. The surface $n^{+}$ electrode of a $p$-type HPGe detector is typically fabricated through thermal diffusion of lithium, resulting in a heavily doped region with a thickness of ~ 0.5–1 mm. However, this process introduces a layer in the $n^{+}$ electrode that is almost inactive close to the surface, with its charge collection efficiency increasing as a function of depth: due to the weak electric field in this region and a higher trapping (or recombination) rate, charge carriers are prone to sudden termination, leading to incomplete charge collection martin_determining_2012 ; aguayo_characteristics_2013 ; lehnert_background_2016 ; dai_modeling_2023 . Thus, charge depositions in this layer are either not registered at all or exhibit a measured energy lower than their true energy. It is sometimes called dead layer, inactive layer, or transition layer in the literature. For clarity, we will call it the reduced charge collection (RCC) layer in this paper. These RCC layer events constitute a significant background component in rare-event searches aalseth_maximum_2015 ; bonet_pulse_2024 ; agostini_pulse_2022 ; li_differentiation_2014 ; yang_bulk_2018 ; li_identification_2022 ; zhang_machine_2024 ; wang_bulk-surface_2025 .
@@
We want to study the charge drift of RCC layer events and develop a pulse shape simulation method that can both deepen our understanding of the underlying detector physics and enable a more precise evaluation of the signal efficiencies of pulse shape discrimination techniques aalseth_maximum_2015 ; alvis_multi-site_2019 ; agostini_pulse_2022 ; comellato_topologies_2023 ; bonet_pulse_2024 ; dai_search_2022 ; zhang_searching_2024 . Realistic simulations can also generate clean datasets with known ground truth for developing advanced discrimination techniques zhang_machine_2024 .
@@ 2 Method
With "sensitive region", we refer to the detector bulk volume with full charge collection efficiency (CCE). RCC layer refers to the outside, including the whole $n$-type neutral region and a small part of the $n$-type depleted region, as shown in Figure 1. Following the convention of previous researches aguayo_characteristics_2013 ; gerda_collaboration_characterization_2019 ; dai_modeling_2023 , the depth from the detector surface to the boundary between the sensitive region and the RCC layer is termed the full charge collection depth (FCCD), which is also conventionally regarded as the depth of the RCC layer.
@@ 2.1 Ionized impurity density profile
Ionized impurities refer to impurities with ionized atoms regardless of whether the carriers are depleted, including donors and acceptors. The ionized impurity density profiles govern the electric field and the physical processes within the RCC layer.
The initial acceptor impurity density is established during crystal growth, though external contaminants may enter during fabrication. This results in a complex three-dimensional distribution that is impossible to predict via first-principles modeling and difficult to measure experimentally after processing. Fortunately, this distribution can be inferred through capacitance-voltage measurements hagemann_determination_2024 . In contrast, donor impurities are introduced primarily through lithium thermal diffusion from the $n^{+}$ electrode. The resulting profile is described by the complementary error function (erfc), representing the solution to the diffusion equation for a semi-infinite medium with a constant surface concentration. The saturated surface lithium concentration and the diffusion coefficient are modeled by empirical Arrhenius-type fits taken from the literature sangster_ge-_1997 ; fuller_diffusion_1953 ; fuller_mobility_1954 ; dai_modeling_2023 .
Therefore, the net ionized impurity density profile throughout the crystal is described by
[…]
$N_{\text{d}}(d)=N_{\text{s}}\cdot\text{erfc}(\frac{d}{2\sqrt{D_{\text{Li}}t_{\text{an}}}})\,$, (2)
$N_{\text{s}}=N_{\text{s,0}}\cdot\exp(-\frac{T_{0}}{T_{\text{an}}})\,$, (3)
$D_{\text{Li}}=D_{\text{Li,0}}\cdot\exp(-\frac{H}{RT_{\text{an}}})\,$, (4)
@@ 2.2 Charge-carrier mobilities
Figure 3 presents the electron and hole mobility curves calculated using the parameters described above. When getting closer to the $n^{+}$ surface, the heavily doped lithium impurities significantly reduce the carrier mobility compared to the bulk sensitive region due to impurity scattering.
It is important to note that this work treats the mobility as a low-field quantity and neglects velocity saturation. In high-field regions, such as the area near the groove where the field can reach several kV/cm, this approximation leads to an overestimation of the drift velocity. However, because the total drift time for RCC-layer events is dominated by the slow transport within the heavily doped RCC layer, omitting the field dependence should introduce only a tiny bias to the overall signal shape. A more comprehensive drift model that incorporates both field-dependent mobility and crystal orientation effects is reserved for future work.
@@ 2.3 Charge-carrier diffusion, self-repulsion, and trapping
The trapping effect of carriers can be described by a lifetime model. Ideally, the carrier lifetime should be treated as depth-dependent, varying with the density of trapping centers. In this work, as a first-order approximation, we employ a piecewise constant-lifetime trapping model dai_modeling_2023 ; boggs_numerical_2023 , assume different effective lifetimes in the sensitive region and the RCC layer. Therefore, our trapping model utilizes four parameters: electron/hole lifetimes in the RCC layer $\tau_{\text{e/h,R}}$, and in the sensitive region $\tau_{\text{e/h,S}}$, respectively. While the simplified lifetime profile is a necessary approximation that has also proven effective in Section 4.4, it introduces a potential uncertaintiy that is currently challenging to assess. This limitation is discussed further in Section 4.4. In addition, to mitigate the high computational cost of simulating trapping effects via pure Monte Carlo sampling, this work adopts an analytical approach in which an exponential decay is applied after the charge drift.
For this paper, we set $\tau_{\text{e/h,S}}$ to 1 ms knoll_radiation_2010 , a typical value for high-purity germanium. However, in the RCC layer, the extremely high donor impurity density makes carriers more prone to trapping, resulting in significantly shorter lifetimes than those in the sensitive region. The lifetimes $\tau_{\text{e/h,R}}$ must be determined by experimental calibration. According to the Shockley-Ramo theorem he_review_2001 , the induced signal $Q_{\text{e/h}}(t_{i})$ is given by
$Q_{\text{e/h}}(t_{i})$ $=Q_{\text{e/h, A}}(t_{i})+Q_{\text{e/h, TR}}(t_{i})+Q_{\text{e/h, TS}}(t_{i})\,$, (15)
$Q_{\text{e/h, A}}(t_{i})$ $=w(\widehat{r}_{\text{e/h}}(t_{i}))q_{\text{e/h}}(t_{i})\,$, (16)
$Q_{\text{e/h, TR}}(t_{i})$ $=\sum_{i^{\prime}\in I_{\text{R}}(t_{i-1})}{w(\widehat{r}_{\text{e/h}}(t_{i^{\prime}}))q_{\text{e/h}}(t_{i^{\prime}})\frac{dt_{i^{\prime}}}{\tau_{\text{e/h, R}}}}\,$, (17)
$Q_{\text{e/h, TS}}(t_{i})$ $=\sum_{i^{\prime}\in I_{\text{S}}(t_{i-1})}{w(\widehat{r}_{\text{e/h}}(t_{i^{\prime}}))q_{\text{e/h}}(t_{i^{\prime}})\frac{dt_{i^{\prime}}}{\tau_{\text{e/h, S}}}}\,$, (18)
$q_{\text{e/h}}(t_{i})$ $=q_{\text{e/h}}(t_{0})-q_{\text{e/h, TR}}(t_{i})-q_{\text{e/h, TS}}(t_{i})\,$, (19)
$q_{\text{e/h, TR}}(t_{i})$ $=\sum_{i^{\prime}\in I_{\text{R}}(t_{i-1})}q_{\text{e/h}}(t_{i^{\prime}})\frac{dt_{i^{\prime}}}{\tau_{\text{e/h, R}}}\,$, (20)
$q_{\text{e/h, TS}}(t_{i})$ $=\sum_{i^{\prime}\in I_{\text{S}}(t_{i-1})}q_{\text{e/h}}(t_{i^{\prime}})\frac{dt_{i^{\prime}}}{\tau_{\text{e/h, S}}}\,$, (21)
where $t_{i}$ is the $i$-th timestamp of a drift (starting at $t_{0}=0$), $dt_{i}$ ($t_{i+1}-t_{i}$) is the $i$-th time step, $Q_{\text{e/h, A}}(t_{i})$ denotes the induced signal contributed by alive carriers (at the timestamp $t_{i}$, the same below), $Q_{\text{e/h, TR}}(t_{i})$ and $Q_{\text{e/h, TS}}(t_{i})$ denote the induced signal contributed by carriers trapped in the RCC layer and sensitive region respectively, $q_{\text{e/h}}(t_{i})$ denotes the charges of alive carriers, $q_{\text{e/h, TR}}(t_{i})$ and $q_{\text{e/h, TS}}(t_{i})$ denote the charges of carriers trapped in the RCC layer and sensitive region respectively, $\widehat{r}_{\text{e/h}}(t_{i})$ denotes the carrier drift paths, $w(\widehat{r}_{\text{e/h}}(t_{i}))$ denotes the weighting potentials, and $I_{\text{R}}(t_{i})$ and $I_{\text{S}}(t_{i})$ denote the sets of timestamp indices when the carrier is located in the RCC layer and the sensitive region respectively.
@@ 2.4 Charge collection efficiency
Based on the aforementioned physical models, we can simulate both the paths of carriers and the induced signals of RCC layer events. Furthermore, we can calculate the CCE curve of the RCC layer $f_{\text{CCE}}(x)$ and its uncertainty dai_modeling_2023 ; ma_study_2017 with
$f_{\text{CCE}}(x)$ $=\frac{A_{\text{p}}(x)}{Ne}=\frac{A_{\text{p}}(x)}{\frac{E}{E_{0}}e}\,$, (22)
$\frac{\sigma_{f}}{f}$ $=\frac{1}{\sqrt{fN}}\,$, (23)
where $x$ is the depth of the simulated single-site event from the detector surface, $A_{\text{p}}$ is the amplitude of the simulated charge pulse, $E$ is the energy deposited, $N$ is the initial number of the simulated charge carrier pairs, and $E_{0}$ is the ionization energy for germanium. It should be noted that the uncertainty modeled here is overestimated, as the trapping effect is treated analytically after the drift.
The CCE curve is a fundamental property of $p$-type germanium detectors, mainly governed by the impurity density profiles and detector geometry. The applied bias voltage has only a minor impact on the CCE near the depletion boundary. It has been proven to be crucial for interpreting experimental spectra and correcting distortions in the raw simulated energy spectra in the low-energy region because alpha, beta, and low-energy gamma particles from outside have a higher probability to deposit energy close to the surface ma_study_2017 ; dai_modeling_2023 ; aalseth_maximum_2015 ; agostini_pulse_2022 ; bonet_pulse_2024 .
For each event, we denote the simulated deposited energies and depths with $(E_{i},x_{i})$. The raw simulated total energy $E_{\text{raw}}$, the total energy corrected with CCE curve $E_{\text{w/}}$, and total energy corrected without CCE curve $E_{\text{w/o}}$ are given by
@@
Figure 4: Schematic illustration of the CCE curve $f_{\text{CCE}}(x)$ and the step function $\Theta(x-\text{FCCD})$ with FCCD = 0.9 mm.
For $p$-type HPGe detectors, in most cases, electrons in the RCC layer contribute negligibly to the induced signal on the $p^{+}$ contact because the weighting potential there is nearly zero li_identification_2022 ; dai_pulse_2024 ; zhang_searching_2023 . Thus, the CCE approximately equals the fraction of holes that escape the RCC layer before getting trapped or recombined, reach the sensitive region, and get collected by the $p^{+}$ electrode. One can determine the FCCD through pulse shape simulation but also experimentally, by matching the experimentally measured energy spectrum with the energy spectrum simulated using a Monte-Carlo software like Geant4 jiang_measurement_2016 ; ma_study_2017 ; aguayo_characteristics_2013 ; gerda_collaboration_characterization_2019 ; agostinelli_geant4simulation_2003 ; allison_geant4_2006 ; allison_recent_2016 .
@@ 3 Method validation
In the theory-based approach, we compute the CCE in the RCC layer by solving the one-dimensional hole transport equation analytically. For this, we consider three hole-transport physics processes: drift under the electric field, diffusion, and trapping.
@@ 3.2 Charge collection efficiency validation
Figure 7 shows the distribution of the time when the holes arrive at the $p$-$n$ boundary for a starting depth of 0.5 mm from the surface. The integral of this curve divided by the initial hole number (3389) is the charge collection efficiency of the event, see Equation 22. The simulated distribution agrees with the analytical one. Notably, the uncertainties in the simulated histogram are overestimated, similar to the calculation of CCE uncertainty described in Section 2.4.
@@
In addition to the arrival time, we compare the depth dependency of the CCE by generating events at different depths in the RCC layer. Figure 8 shows that the two CCE curves from the numerical simulation and analytical solution are also consistent.